# Rationale-Enhanced Decoding for Multi-modal Chain-of-Thought

### v1

@@ 3.2 Rationale-Enhanced Decoding (RED)
where $Z_{\theta}=\sum_{w\in\mathcal{V}}p_{\theta}(y_{i}=w|\bm{y}_{<i},x,q)\times p_{\theta}(y_{i}=w|\bm{y}_{<i},r,q)^{\lambda}$ is the normalization constant, and $\lambda$ is a hyperparameter for modulating the influence of $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$. Intuitively, Eq. (8) is a power-of-experts Hinton2002_poe emphasizing overlap between image- and rationale conditional probabilities. Indeed, Eq. (8) is the closed-form solution maximizing Eq. (7) as follows.
Proposition 1.
@@
Consider the KL-constrained reward maximization objective Schulman_ICML15_TRPO ; Schulman_ICML15_PPO ; Rafailov_NeurIPS23_DPO :
@@
where $R$ is a reward function, $s$ is a state (input context), and $a$ is an action (output). Obviously, Eq. (7) is a special case of Eq. (9). From Appendix A.1 of Rafailov_NeurIPS23_DPO , the optimal policy $\pi^{*}(a|s)$ for this objective is given by
@@
where $Z(s)=\sum_{a^{\prime}}\pi_{\mathrm{ref}}(a^{\prime}|s)\exp(\frac{1}{\beta}R(s,a^{\prime}))$ is the partition function. In the maximization of Eq. (7), given $a=y_{i}$ and $s=(\bm{y}_{i<},x,q)$, we can set each component in Eq. (9) as $\pi_{\mathrm{ref}}(a|s)=p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$, $R(s,a)=\log p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, and $\beta=1/\lambda$. Substituting them into the optimal policy of Eq. (10) yields
@@
Proposition 1 shows that sampling by Eq. (8) yields the optimal distribution maximizing $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ at decoding time without any auxiliary reward models. This theoretical property supports the reliability of RED as a method for improving LVLMs in the multi-modal CoT. We also show a comparison to other possible alternatives to validate the efficacy of Eq. (8) in Appendix B.
@@ 3.3 Algorithm
$\hat{\operatorname{logits}}_{\theta}(y_{i}):=\log\operatorname{softmax}(\operatorname{logits}_{\theta}(y_{i}|\bm{y}_{<i},x,q))+\lambda\log\operatorname{softmax}(\operatorname{logits}_{\theta}(y_{i}|\bm{y}_{<i},r,q)).$ (12)
@@
$\hat{p}_{\theta}(y_{i})$ $=\exp\log(\hat{p}_{\theta}(y_{i}))=\exp(\log(p_{\theta}(y_{i}|\bm{y}_{<i},x,q))+\log(p_{\theta}(y_{i}|\bm{y}_{<i},r,q)^{\lambda})-\log Z_{\theta})$
$\propto\frac{\exp(\log(p_{\theta}(y_{i}|\bm{y}_{<i},x,q)\times p_{\theta}(y_{i}|\bm{y}_{<i},r,q)^{\lambda}))}{\sum_{{w}\in\mathcal{V}}\exp(\log(p_{\theta}(y_{i}=w|\bm{y}_{<i},x,q)\times p_{\theta}(y_{i}=w|\bm{y}_{<i},r,q)^{\lambda}))}.$ (13)
Thus, we can implement RED by computing the weighted sum of the log-softmax logits as the new logits for the next token prediction. We show the overall procedures of RED in Algorithm 1.
@@
1:  $r\leftarrow\operatorname{generate}_{\theta}(x,q)$    # Generate arbitrary rationales
2:  $\bm{y}\leftarrow\emptyset$
@@ 4 Experiments
We validate the efficacy of RED via comprehensive evaluation on multiple multi-modal reasoning datasets with pre-trained LVLMs, comparing to existing multi-modal CoT methods and plug-and-play decoding baselines. We also analyze interventions on rationale quality.
4.1 Settings
Baselines. The baselines include: Baseline (standard inference with $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$), CoT Wei_NeurIPS22_CoT ; Zhang_TMLR_mm-cot (prompting to generate text rationales), and CCoT Mitra_CVPR24_CCoT (a state-of-the-art training-free method generating JSON-formatted scene graph rationales). Other plug-and-play decoding baselines are: VCD Leng_CVPR24_VCD , contrasting $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ with $p_{\theta}(y_{i}|\bm{y}_{<i},x^{\prime},q)$, where $x^{\prime}$ is a corrupted input image by adding Gaussian noise, and VCD + ICD Wang_ACL24F_ICD , improving VCD using a variant query $q^{\prime}$ modified by adversarial instruction. While they were proposed for object hallucination in LVLMs, comparing RED with them helps validate the practicality because they reportedly improved general performance Leng_CVPR24_VCD ; Wang_ACL24F_ICD .
Benchmark Datasets. We used six diverse visual reasoning benchmark datasets of various domains: GQA Hudson_CVPR19_gqa , TextVQA Singh_CVPR19_textvqa , MME Fu_2023_mme , SEED-I Li_CVPR24_seed_bench , LLaVA-Bench Liu_CVPR24_improved_llava_llavabench , and MM-Vet Yu_ICML24_mmvet .
Models. We used publicly available LVLMs on HuggingFace Wolf_arXiv19_huggingface : Gemma-3 (4B, 12B, 27B) Google_2025_gemma3 , Qwen-2.5-VL (7B, 32B, 72B) Bai_2025_Qwen-vl , and Llama3-LLaVA-Next (8B) Li_2024_llavanext .
Evaluation Protocols. We used greedy decoding. $\lambda$ for RED was chosen from {0.1, 0.3, 0.5, 1.0, 10.0} using scores on validation or development sets; Appendix B.3 discusses the effect of $\lambda$. Tables show scores with parenthesized relative improvement or degradation from Baseline.
4.2 Evaluation on General Visual Reasoning Tasks
Table 2 shows RED’s general visual reasoning improvements. While CoT/CCoT sometimes outperformed Baseline, they were inconsistent, with large drops on some LVLM-benchmark pairs, especially TextVQA, which requires understanding textual content in images. In contrast, RED consistently improved CoT/CCoT and outperformed all baselines in nearly all cases, even on TextVQA, where others degraded significantly. RED thus successfully addresses the issues of the existing multi-modal CoT using $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ in terms of performance. The consistent improvements offered by RED highlight the potential value of intermediate rationales generated in multi-modal CoT and its practicality for leveraging in diverse domains.
@@ 4.2 Evaluation on General Visual Reasoning Tasks
GQA LLaVA-Bench MME MMVet SEED-I TextVQA
@@
Baseline 40.00 73.20 1211.34 370.00 44.00 65.39 63.95
VCD 38.74 (-1.26) 75.70 (+2.50) 1184.11 (-27.23) 338.57 (-31.43) 44.80 (+0.80) 65.19 (-0.20) 64.02 (+0.07)
VCD + ICD 38.62 (-1.33) 73.30 (+0.10) 1182.29 (-29.05) 340.71 (-29.29) 46.70 (+2.70) 65.20 (-0.19) 64.03 (+0.08)
CoT 41.08 (+1.08) 72.90 (-0.30) 1254.77 (+43.43) 341.07 (-28.93) 46.20 (+2.20) 65.90 (+0.51) 60.62 (-3.33)
CCoT 44.54 (+4.54) 73.60 (+0.40) 1294.46 (+83.12) 396.43 (+26.43) 46.50 (+2.50) 66.63 (+1.24) 63.23 (-0.72)
CoT + RED 42.19 (+2.19) 76.30 (+3.10) 1325.77 (+114.43) 645.35 (+275.35) 46.60 (+2.60) 66.89 (+1.50) 65.13 (+1.18)
CCoT + RED 45.87 (+5.87) 76.90 (+3.70) 1330.07 (+118.73) 611.43 (+241.43) 49.10 (+5.10) 66.73 (+1.34) 65.54 (+1.59)
@@
Baseline 45.34 79.00 1171.37 545.71 57.70 71.01 69.81
VCD 44.11 (-1.23) 78.40 (-0.60) 1152.02 (-19.34) 543.93 (-1.79) 58.30 (+0.60) 70.96 (-0.05) 69.40 (-0.51)
VCD + ICD 44.16 (-1.18) 79.70 (+0.70) 1153.86 (-17.50) 651.79 (+106.07) 57.40 (-0.30) 71.01 (+0.00) 69.50 (-0.41)
CoT 41.76 (-3.58) 78.90 (-0.10) 1507.67 (+336.30) 661.07 (+115.36) 57.50 (-0.20) 70.75 (-0.26) 66.23 (-3.58)
CCoT 44.50 (-0.84) 79.00 (+0.00) 1289.67 (+118.30) 604.29 (+58.58) 53.00 (-4.70) 71.69 (+0.68) 69.70 (-0.11)
CoT + RED 46.07 (+0.73) 81.00 (+2.00) 1574.52 (+403.15) 695.36 (+149.65) 59.50 (+1.80) 72.15 (+1.14) 70.73 (+0.92)
CCoT + RED 47.50 (+2.16) 80.60 (+1.60) 1359.28 (+187.91) 642.50 (+96.79) 58.40 (+0.70) 72.76 (+1.75) 71.09 (+1.28)
@@
Baseline 60.88 82.10 1665.22 621.07 56.70 58.13 77.76
VCD 59.34 (-1.54) 82.20 (+0.10) 1650.80 (-34.42) 631.07 (+10.00) 58.30 (-0.40) 60.32 (+2.19) 77.53 (-0.23)
VCD + ICD 59.40 (-1.48) 82.40 (+0.30) 1584.17 (-101.04) 651.79 (+30.71) 57.50 (-1.20) 60.71 (+2.58) 75.45 (-2.31)
CoT 46.70 (-14.18) 81.80 (-0.30) 1555.52 (-109.70) 705.00 (+83.93) 56.80 (+0.10) 60.49 (+2.36) 70.32 (-7.44)
CCoT 46.69 (-14.19) 81.00 (-1.10) 1559.71 (-105.51) 634.29 (+13.22) 52.90 (-3.80) 73.18 (+15.05) 74.60 (-3.16)
CoT + RED 61.06 (+0.18) 82.20 (+0.10) 1706.25 (+41.03) 706.79 (+86.72) 60.70 (+4.00) 76.50 (+18.37) 77.98 (+0.22)
CCoT + RED 61.92 (+1.04) 84.60 (+2.50) 1704.83 (+39.61) 648.21 (+27.14) 56.80 (+0.10) 78.37 (+20.24) 78.61 (+0.85)
@@
Baseline 65.22 65.60 1583.67 332.14 37.70 72.57 65.01
VCD 63.60 (-1.62) 67.60 (+2.00) 1512.10 (-71.57) 341.79 (+9.64) 40.50 (+2.80) 71.93 (-0.64) 63.51 (-1.50)
VCD + ICD 63.88 (-1.34) 70.50 (+4.90) 1485.12 (-98.55) 317.50 (+14.64) 40.30 (+2.60) 71.71 (-0.86) 62.97 (-2.03)
CoT 61.82 (-3.40) 65.60 (+0.00) 1462.63 (-121.04) 415.71 (+83.57) 42.10 (+4.40) 72.38 (-0.19) 64.05 (-0.96)
CCoT 60.43 (-4.79) 63.50 (-2.10) 1424.86 (-158.81) 343.57 (+11.43) 35.80 (-1.90) 72.74 (+0.17) 64.18 (-0.83)
CoT + RED 65.48 (+0.26) 74.10 (+8.50) 1583.56 (-0.11) 410.36 (+78.22) 40.60 (+2.90) 73.19 (+0.62) 66.66 (+1.65)
CCoT + RED 65.91 (+0.69) 76.50 (+10.90) 1562.72 (-20.95) 441.79 (+109.65) 42.10 (+4.40) 73.22 (+0.65) 66.04 (+1.03)
4.3 Detailed Analysis for Multi-modal Reasoning Capabilities
Table 3: Detailed performance analysis on SEED-I for different LVLMs and CoT strategies.

### v2

@@ 4.2 Rationale-Enhanced Decoding (RED)
where $Z_{\theta}=\sum_{w\in\mathcal{V}}p_{\theta}(y_{i}=w|\bm{y}_{<i},x,q)\times p_{\theta}(y_{i}=w|\bm{y}_{<i},r,q)^{\lambda}$ is the normalization constant, and $\lambda$ is a hyperparameter for modulating the influence of $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$. Intuitively, Eq. (8) is a power-of-experts [17] emphasizing overlap between image- and rationale conditional probabilities. Indeed, Eq. (8) is the closed-form solution maximizing Eq. (7) as follows.
Theorem 4.1.
@@
Consider the KL-constrained reward maximization objective [46, 47, 44]:
@@
where $R$ is a reward function, $s$ is a state (input context), and $a$ is an action (output). Obviously, Eq. (7) is a special case of Eq. (9). From Appendix A.1 of [44], the optimal policy $\pi^{*}(a|s)$ for this objective is given by
@@
where $Z(s)=\sum_{a^{\prime}}\pi_{\mathrm{ref}}(a^{\prime}|s)\exp(\frac{1}{\beta}R(s,a^{\prime}))$ is the partition function. In the maximization of Eq. (7), given $a=y_{i}$ and $s=(\bm{y}_{i<},x,r,q)$, we can set each component in Eq. (9) as $\pi_{\mathrm{ref}}(a|s)=p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$, $R(s,a)=\log p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, and $\beta=1/\lambda$. Substituting them into the optimal policy of Eq. (10) yields
@@
Theorem 4.1 shows that sampling by Eq. (8) yields the optimal distribution maximizing $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ at decoding time without any auxiliary reward models. This theoretical property supports the reliability of RED as a method for improving LVLMs in the multi-modal CoT. We also show a comparison to other possible alternatives to validate the efficacy of Eq. (8) in Appendix.
@@ 4.3 Algorithm
$\hat{\operatorname{logits}}_{\theta}(y_{i}):=\log\operatorname{softmax}(\operatorname{logits}_{\theta}(y_{i}|\bm{y}_{<i},x,q))$
$+\lambda\log\operatorname{softmax}(\operatorname{logits}_{\theta}(y_{i}|\bm{y}_{<i},r,q)).$ (12)
@@
$\hat{p}_{\theta}(y_{i})=\exp\log(\hat{p}_{\theta}(y_{i}))$
$=\exp(\log(p_{\theta}(y_{i}|\bm{y}_{<i},x,q))\!+\!\log(p_{\theta}(y_{i}|\bm{y}_{<i},r,q)^{\lambda})\!-\!\log Z_{\theta})$
$\propto\!\operatorname{softmax}(\log(p_{\theta}(y_{i}|\bm{y}_{<i},\!x,\!q)\!\times\!p_{\theta}(y_{i}|\bm{y}_{<i},\!r,\!q)^{\lambda})).$ (13)
Thus, we can implement RED by computing the weighted sum of the log-softmax logits as the new logits for the next token prediction. To avoid the latency overhead, we simultaneously compute $\operatorname{logits}_{\theta}(y|\bm{y},x,q)$ and $\operatorname{logits}_{\theta}(y|\bm{y},r,q)$ by batch parallel inference. We show the overall procedures of RED in Algorithm 1.
@@
[…]
@@ 5 Experiments
We validate the efficacy of RED via comprehensive evaluation on multiple multi-modal reasoning datasets with pre-trained LVLMs, comparing to existing multi-modal CoT methods and plug-and-play decoding baselines. We also analyze interventions on the rationale quality.
@@
GQA LLaVA-Bench MME MMVet SEED-I TextVQA MathVista
@@
Baseline 40.00 73.20 1211.34 370.00 44.00 65.39 63.95 41.40
VCD 38.74 (-1.26) 75.70 (+2.50) 1184.11 (-27.23) 338.57 (-31.43) 44.80 (+0.80) 65.19 (-0.20) 64.02 (+0.07) 39.30 (-2.10)
VCD + ICD 38.62 (-1.33) 73.30 (+0.10) 1182.29 (-29.05) 340.71 (-29.29) 46.70 (+2.70) 65.20 (-0.19) 64.03 (+0.08) 40.40 (-1.00)
CoT 41.08 (+1.08) 72.90 (-0.30) 1254.77 (+43.43) 341.07 (-28.93) 46.20 (+2.20) 65.90 (+0.51) 60.62 (-3.33) 40.10 (-1.30)
CCoT 44.54 (+4.54) 73.60 (+0.40) 1294.46 (+83.12) 396.43 (+26.43) 46.50 (+2.50) 66.63 (+1.24) 63.23 (-0.72) 41.10 (-0.30)
CoT + RED 42.19 (+2.19) 76.30 (+3.10) 1325.77 (+114.43) 645.35 (+275.35) 46.60 (+2.60) 66.89 (+1.50) 65.13 (+1.18) 43.50 (+2.10)
CCoT + RED 45.87 (+5.87) 76.90 (+3.70) 1330.07 (+118.73) 611.43 (+241.43) 49.10 (+5.10) 66.73 (+1.34) 65.54 (+1.59) 42.00 (+0.60)
@@
Baseline 45.34 79.00 1171.37 545.71 57.70 71.01 69.81 52.10
VCD 44.11 (-1.23) 78.40 (-0.60) 1152.02 (-19.34) 543.93 (-1.79) 58.30 (+0.60) 70.96 (-0.05) 69.40 (-0.51) 51.50 (-0.60)
VCD + ICD 44.16 (-1.18) 79.70 (+0.70) 1153.86 (-17.50) 651.79 (+106.07) 57.40 (-0.30) 71.01 (+0.00) 69.50 (-0.41) 51.80 (-0.30)
CoT 41.76 (-3.58) 78.90 (-0.10) 1507.67 (+336.30) 661.07 (+115.36) 57.50 (-0.20) 70.75 (-0.26) 66.23 (-3.58) 53.50 (+1.40)
CCoT 44.50 (-0.84) 79.00 (+0.00) 1289.67 (+118.30) 604.29 (+58.58) 53.00 (-4.70) 71.69 (+0.68) 69.70 (-0.11) 51.20 (-0.90)
CoT + RED 46.07 (+0.73) 81.00 (+2.00) 1574.52 (+403.15) 695.36 (+149.65) 59.50 (+1.80) 72.15 (+1.14) 70.73 (+0.92) 54.80 (+2.70)
CCoT + RED 47.50 (+2.16) 80.60 (+1.60) 1359.28 (+187.91) 642.50 (+96.79) 58.40 (+0.70) 72.76 (+1.75) 71.09 (+1.28) 53.50 (+1.40)
@@
Baseline 60.88 82.10 1665.22 621.07 56.70 58.13 77.76 64.70
VCD 59.34 (-1.54) 82.20 (+0.10) 1650.80 (-34.42) 631.07 (+10.00) 58.30 (-0.40) 60.32 (+2.19) 77.53 (-0.23) 64.50 (-0.20)
VCD + ICD 59.40 (-1.48) 82.40 (+0.30) 1584.17 (-101.04) 651.79 (+30.71) 57.50 (-1.20) 60.71 (+2.58) 75.45 (-2.31) 64.80(+0.10)
CoT 46.70 (-14.18) 81.80 (-0.30) 1555.52 (-109.70) 705.00 (+83.93) 56.80 (+0.10) 60.49 (+2.36) 70.32 (-7.44) 61.30 (-3.40)
CCoT 46.69 (-14.19) 81.00 (-1.10) 1559.71 (-105.51) 634.29 (+13.22) 52.90 (-3.80) 73.18 (+15.05) 74.60 (-3.16) 61.30 (-3.40)
CoT + RED 61.06 (+0.18) 82.20 (+0.10) 1706.25 (+41.03) 706.79 (+86.72) 60.70 (+4.00) 76.50 (+18.37) 77.98 (+0.22) 70.60 (+5.90)
CCoT + RED 61.92 (+1.04) 84.60 (+2.50) 1704.83 (+39.61) 648.21 (+27.14) 56.80 (+0.10) 78.37 (+20.24) 78.61 (+0.85) 68.10 (+3.40)
@@
Baseline 65.22 65.60 1583.67 332.14 37.70 72.57 65.01 37.70
VCD 63.60 (-1.62) 67.60 (+2.00) 1512.10 (-71.57) 341.79 (+9.64) 40.50 (+2.80) 71.93 (-0.64) 63.51 (-1.50) 35.90 (-1.80)
VCD + ICD 63.88 (-1.34) 70.50 (+4.90) 1485.12 (-98.55) 317.50 (+14.64) 40.30 (+2.60) 71.71 (-0.86) 62.97 (-2.03) 36.20 (-1.50)
CoT 61.82 (-3.40) 65.60 (+0.00) 1462.63 (-121.04) 415.71 (+83.57) 42.10 (+4.40) 72.38 (-0.19) 64.05 (-0.96) 36.50 (-1.20)
CCoT 60.43 (-4.79) 63.50 (-2.10) 1424.86 (-158.81) 343.57 (+11.43) 35.80 (-1.90) 72.74 (+0.17) 64.18 (-0.83) 36.70 (-1.00)
CoT + RED 65.48 (+0.26) 74.10 (+8.50) 1583.56 (-0.11) 410.36 (+78.22) 40.60 (+2.90) 73.19 (+0.62) 66.66 (+1.65) 39.10 (+1.40)
CCoT + RED 65.91 (+0.69) 76.50 (+10.90) 1562.72 (-20.95) 441.79 (+109.65) 42.10 (+4.40) 73.22 (+0.65) 66.04 (+1.03) 38.80 (+1.10)
Table 3: Detailed performance analysis on SEED-I for different CoT strategies. RED’s outputs evolve in accordance with the characteristics of specified CoT strategies. For example, CoT with natural text rationales enhances text understanding, while CCoT with structured scene graph rationales improves visual reasoning capability.
@@
5.1 Settings
Baselines. The baselines include: Baseline (standard inference with $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$), CoT [60, 71] (prompting to generate text rationales), and CCoT [37] (a state-of-the-art training-free method generating JSON-formatted scene graph rationales). Other plug-and-play decoding baselines are: VCD [26], contrasting $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ with $p_{\theta}(y_{i}|\bm{y}_{<i},x^{\prime},q)$, where $x^{\prime}$ is a corrupted input image by adding Gaussian noise, and VCD + ICD [58], improving VCD using a variant query $q^{\prime}$ modified by adversarial instruction. While they were proposed for object hallucination in LVLMs, comparing RED with them helps validate the practicality because they reportedly improved general performance [26, 58].
Benchmark Datasets. We used six diverse visual reasoning benchmark datasets: GQA [18], TextVQA [49], MME [11], SEED-I [27], LLaVA-Bench [33], MM-Vet [65], and MathVista [35]. These benchmarks are generally used for evaluating various types of multi-modal capabilities, including general question answering, visual recognition, OCR, mathematical reasoning, multi-modal comprehension, and spatial understanding.
Models. We used publicly available LVLMs on HuggingFace [61]: Gemma-3 (4B, 12B, 27B) [55], Qwen-2.5-VL (7B, 32B, 72B) [3], and Llama3-LLaVA-Next (8B) [28].
Evaluation Protocols. We used greedy decoding. $\lambda$ for RED was chosen from {0.1, 0.3, 0.5, 1.0, 10.0} using scores on validation or development sets; We discuss the effect of $\lambda$ in Appendix. Tables show scores with parenthesized relative improvement or degradation from Baseline.
5.2 Evaluation on General Visual Reasoning Tasks
Table 2 shows RED’s general visual reasoning improvements. While CoT/CCoT sometimes outperformed Baseline, they were inconsistent, with large drops on some LVLM-benchmark pairs, especially TextVQA, which requires understanding texts in images. In contrast, RED consistently improved CoT/CCoT and outperformed all baselines in nearly all cases, even on TextVQA. RED thus successfully addresses the issues of the existing multi-modal CoT using $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ in terms of performance. The consistent improvements offered by RED highlight the potential value of intermediate rationales generated in multi-modal CoT and its practicality for leveraging in diverse domains.
5.3 Detailed Analysis for Multi-modal Capabilities